# Densities Of Primes And Primitive Roots

### v13

@@ Abstract
Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)=d\>|\>p-1$ be the order of $u\operatorname{mod}p$. This note provides an effective lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-1}$ for the number of primes $p\leq x$ with a fixed primitive root $u\text{ mod }p$ for all large numbers $x\geq 1$. The current results in the literature have the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$, and restrictions on the fixed primitive root to a subset of at least three or more integers. November 15, 2025
@@ 1 Introduction
The symbol $\mathbb{Z}=\{0,\pm 1,\pm 2,\ldots\}$ denotes the set of integers, and the symbol $\mathbb{P}=\{2,3,5,\ldots\}$ denotes the set of prime numbers. The constant $\delta(u)\geq 0$ is the density of the subset of primes
$\mathcal{P}_{u}=\left\{p\in\mathbb{P}:\operatorname{ord}_{p}(u)=p-1\right\}\subset\mathbb{P}$ (1.1)
with a fixed primitive root $u\in\mathbb{Z}$. Let $u\neq\pm 1,v^{2}$ be a fixed integer, and let $x\geq 1$ be a large number. The expected number of primes $p\leq x$ with a fixed primitive root $u\operatorname{mod}p$ has the asymptotic formula
@@
A conditional proof of this result was achieved in [25], and simplified sketches of the proof appear in [32, p. 8], and similar references. The determination of the constant $\delta(u)\geq 0$ for a fixed integer $u\in\mathbb{Z}$ is an interesting technical problem, [25, p. 218], [29], [30], et alii. An introduction to its historical development, and its calculations is covered in [32, pp. 3–10], and [44].
@@
$x^{-1}\sum_{u\leq x}\pi_{u}(x)=a_{0}\operatorname{li}(x)+O\left(x(\log x)^{-B}\right),$ (1.3)
where $a_{0}=\prod_{p\geq 2}\left(1-p^{-1}(p-1)^{-1}\right)$ is Artin constant, and $B>1$ is an arbitrary number, was proved in [20] unconditionally, and refined in [44]. These works had shown that almost all admissible integers $u\in\mathbb{Z}-\left\{-1,1,v^{2}:v\in\mathbb{Z}\right\}$ are primitive roots for infinitely many primes. The number of exceptions is a subset of zero density in $\mathbb{Z}$. The individual quantity $\pi_{u}(x)$ in (1.2) can be slightly different from the average quantity in (1.3). The variations, discovered by the Lehmers using numerical experiments, depend on the primes decomposition of the fixed value $u$, see [25, p. 220], [32, p. 3], and similar references for the exact formula for the density $\delta(u)\geq 0$.
The Artin primitive root conjecture for functions fields was proved by Bilharz, see [42, Theorem 10.11], and the same conjecture for polynomials over finite fields was proved in [40]. However, there is no known infinite sequence of primes with a fixed primitive root. The current literature has results infinite sequences of primes with unknown primitive root in a small finite set. For example, in [21] it was proved that for a fixed primes triple $q>r>s\geq 2$, the subset of integers
$\mathcal{A}(q,r,s)=\left\{q^{a}r^{b}s^{c}:0\leq a,b,c\leq 3\right\},$ (1.4)
contains a primitive root for infinitely many primes. This result was later reduced to the smaller subset
$\mathcal{B}(q,r,s)=\left\{q^{a}r^{b}s^{c}:0\leq a,b,c\leq 1\right\},$ (1.5)
see [24]. In both of these results, the lower bound for the number of primes $p\leq x$ with a fixed primitive root in either of the subset $\mathcal{A}(q,r,s)$ or $\mathcal{B}(q,r,s)$ has the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$ for all large number $x\geq 1$.
These results have been extended to quadratic numbers fields in [8], [38], [43], and most recently in [1]. Other related results are given in [33], [18], [39], et alii.
The technique explored in this note provides an improved lower bound for the number of primes $p\leq x$ with a fixed primitive root $u\operatorname{mod}p$ for all large number $x\geq 1$. And the fixed primitive root $u$ is not restricted to a small finite subset such as $\mathcal{A}(q,r,s)$ or $\mathcal{B}(q,r,s)$.
@@
where $\operatorname{li}(x)$ is the logarithm integral, and $\delta(u)>0$ is a constant, for all large numbers $x\geq 1$.
Section 2 to Section 7 introduce the notation and some standard results related to or applicable to the investigation of primitive roots in cyclic groups. The last section presents a proof of Theorem 1.1.
2 Exponentials and Character Sums
Afew standard definitions and other basic results in the theory of exponential and character sums are reviewed in this section. All the estimates are unconditional.
2.1 Simple Characters Sums
Let $G$ be a finite group of order q=#Gq=\#G. The order $\operatorname{ord}(u)$ of an element $u\in G$ is the smallest integer $d\mid q$ such that $u^{d}=1$. An element $\tau\in G$ is called a primitive element if it has order $\operatorname{ord}(\tau)=q$. A cyclic group $G$ is a group generated by a primitive element $\tau\in G$. Given a primitive root $\tau\in G$, every element $0\neq u\in G$ in a cyclic group has a representation as $u=\tau^{v},0\leq v<q$. The integer $v=\log u$ is called the discrete logarithm of $u\neq 0$ with respect to $\tau$.
A character $\chi$ modulo $q\geq 2$, is a complex-valued periodic function $\chi:\mathbb{N}\longrightarrow\mathbb{C}$, and it has order $\operatorname{ord}(\chi)=d\geqslant 1$ if and only if $\chi$$(n)^{d}=1$ for all integers $n\in\mathbb{N},\gcd(n,q)=1$. For $q\neq 2^{r},r\geq 2$, a multiplicative character $\chi$ of order $\operatorname{ord}(\chi)=d\mid q$ has a representation as
$\chi(u)=e^{i2\pi k\log u/(p-1)},$ (2.1)
where $v=\log u$ is the discrete logarithm of $u\neq 0$ with respect to some primitive root, and for some $k\geq 1$, see [31, p. 187], [36, p. 118], and [27, p. 271] for more details.
@@ 2.1 Simple Characters Sums
For a fixed integer $u\neq 0$, and an integer $q\in\mathbb{N}$, let $\chi\neq 1$ be nonprincipal character mod $q$, then
@@
An additive character $\psi$ of order $\operatorname{ord}(\psi)=q$ has a representation as
$\psi(n)=e^{i2\pi kn/q},$ (2.2)
for some $k\geq 1,\gcd(k,q)=1$, see [31, p. 187], [36, p. 118], and [27, p. 271]. The additive character sums are quite similar to Lemma 2.1.
@@
For a fixed integer $u$, and an integer $q\in\mathbb{N}$, let $\psi$ be an additive character of order $\operatorname{ord}\psi=q$, then
@@
3 Representations of the Characteristic Functions
he characteristic function $\Psi:G\longrightarrow\{0,1\}$ of primitive elements is one of the standard analytic tools employed to investigate the various properties of primitive roots in cyclic groups $G$. Many equivalent representations of the characteristic function $\Psi$ of primitive elements are possible. Several of these representations are studied in this section.
3.1 Primitive Roots Tests
For a prime $p\geq 2$, the multiplicative group of the finite fields $\mathbb{F}_{p}$ is a cyclic group for all primes.
Definition 3.1.
The order $\min\{k\in\mathbb{N}:u^{k}\equiv 1\bmod p\}$ of an element $u\in\mathbb{F}_{p}$ is denoted by $\operatorname{ord}_{p}(u)$. An element is a primitive root if and only if $\operatorname{ord}_{p}(u)=p-1$.
Lemma 3.1.
Let $u\in\mathbb{F}_{p}$. Then, $u$ is a primitive ro0t if and only if $u^{(p-1)/q}\not\equiv 1\bmod p$ for all prime divisors $q\mid p-1$.
Basically, this is the classical Lucas-Lehmer primitive root test on the cyclic group $\mathbb{F}_{p}$.
3.2 Divisors Dependent Characteristic Function
A representation of the characteristic function dependent on the orders of the cyclic groups is given below. This representation is sensitive to the primes decompositions $q=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{t}^{e_{t}}$, with $p_{i}$ prime and $e_{i}\geq 1$, of the orders of the cyclic groups q=#Gq=\#G.
Definition 3.2.
The order of an element in the cyclic group $\mathbb{F}_{p}^{\times}$ is defined by $\operatorname{ord}_{p}(v)=\min\{k:v^{k}\equiv 1\bmod p\}$. Primitive elements in this cyclic group have order p1=#Gp-1=\#G.
Lemma 3.2.
Let $G$ be a finite cyclic group of order p1=#Gp-1=\#G, and let $0\neq u\in G$ be an invertible element of the group. Then
[…]
The works in [9], and [48] attribute this formula to Vinogradov. The proof and other details on the characteristic function are given in [14, p. 863], [31, p. 258], [33, p. 18]. The characteristic function for multiple primitive roots is used in [7, p. 146] to study consecutive primitive roots. In [10], [28] it is used to study the gap between primitive roots with respect to the Hamming metric, et cetera. And in [48] it is used to prove the existence of primitive roots in certain small subsets $A\subset\mathbb{F}_{p}$. In [9] it is used to prove that some finite fields do not have primitive roots of the form $a\tau+b$, with $\tau$ primitive and $a,b\in\mathbb{F}_{p}$ constants. In addition, the Artin primitive root conjecture for polynomials over finite fields was proved in [40] using this formula.
3.3 Divisors Free Characteristic Function
It often difficult to derive any meaningful result using the usual divisors dependent characteristic function of primitive elements given in Lemma 3.2. This difficulty is due to the large number of terms that can be generated by the divisors, for example, $d\mid p-1$, involved in the calculations, see [14], [10] for typical applications and [32, p. 19] for a discussion.
A new divisors-free representation of the characteristic function of primitive element is developed here. This representation can overcomes some of the limitations of its counterpart in certain applications. The divisors representation of the characteristic function of primitive roots, Lemma 3.2, detects the order $\operatorname{ord}_{p}(u)$ of the element $u\in\mathbb{F}_{p}$ by means of the divisors of the totient p1p-1. In contrast, the divisors-free representation of the characteristic function, Lemma 3.3, detects the order $\operatorname{ord}_{p}(u)\geq 1$ of the element $u\in\mathbb{F}_{p}$ by means of the solutions of the equation $\tau^{n}-u=0$ in $\mathbb{F}_{p}$, where $u,\tau$ are constants, and $1\leq n<p-1,\gcd(n,p-1)=1,$ is a variable. Two versions are given: a multiplicative version, and an additive version.
Lemma 3.3.
Let $p\geq 2$ be a prime, and let $\tau$ be a primitive root mod $p$. For a nonzero element $u\in\mathbb{F}_{p}$, the followings hold:
(i) If $\psi\neq 1$ is a nonprincipal additive character of order $\operatorname{ord}\psi=p$, then
[…]
(ii) If $\chi\neq 1$ is a nonprincipal multiplicative character of order $\operatorname{ord}\chi=p-1$, then
[…]
where $\bar{u}$ is the inverse of $u\bmod p$.
Proof.
(i) Set the additive character $\psi(s)=e^{i2\pi as/p}\in\mathbb{C}$. Now, as the index $n\in\mathscr{R}=\{n<p:\gcd(n,p-1)=1\}$ ranges over the integers relatively prime to $\varphi(p)=p-1$, the element $\tau^{n}\in\mathbb{F}_{p}^{\times}$ ranges over the primitive roots modulo $p$. Accordingly, the equation
$a=\tau^{n}-u=0$ (3.4)
has a unique solution $n\in\mathscr{R}$ if and only if the fixed element $u\in\mathbb{F}_{p}$ is a primitive root. This implies that the inner sum in
[…]
collapses to $\sum_{0\leq s<p}e^{i2\pi as/p}=\sum_{0\leq s<p}1=p$. Otherwise, if the element $u\in\mathbb{F}_{p}$ is not a primitive root, then the equation (3.4) has no solution $n\in\mathscr{R}$, and the inner sum in (3.5) collapses to $\sum_{0\leq s<p}e^{i2\pi as/p}=0$, this follows from the geometric series formula $\sum_{0\leq n\leq N-1}w^{n}=(w^{N}-1)/(w-1)$, where $w=e^{i2\pi a/p}\neq 1$ and $N=p$. This completes the verification. ∎
@@ 5 Estimates of Power Exponential Sums
where $\omega=e^{i2\pi/p}$ and $x\leq p-1$.
5.1 Power Exponential Sum with Relatively Prime Index
@@
Let $p$ be a large prime, and let $f(n)=e^{i2\pi b\tau^{n}/p}$, where $\tau$ is a primitive root modulo $p$. Start with the representation
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}},$ (5.2)
see (5.1). Use the inclusion exclusion principle to rewrite the exponential sum as
Now, observe that the term $t=0$ contributes $\varphi(p-1)/p$, and rearranging it yield
$=$ $\sum_{n\leq p-1}\frac{1}{p}\sum_{1\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d)-\frac{\varphi(p-1)}{p}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p-1)}{p}.$
Taking absolute value, and applying Lemma 4.1, and Lemma 4.2, yield
$\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}\right|$
$\leq$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left|\sum_{1\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right|\cdot\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p-1)}{p}$
$\ll$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(2p^{1/2}\log p\right)\cdot\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p-1)}{p}$
$\ll$ $\frac{1}{p}\left(2p^{1/2}\log p\right)\cdot\sum_{1\leq t\leq p-1}\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p-1)}{p}$
$\ll$ $\frac{1}{p}\left(2p^{1/2}\log p\right)\cdot\left(2p^{1+\delta}\log p\right)$
$\ll$ $p^{1/2+\delta}(\log p)^{2},$
where $\delta>0$ is a small number. ∎
A different approach to this result appears in [17, Theorem 6], and related results are given in [16], [22], [6], and [23, Theorem 1]. The upper bound given in Theorem 5.1 seems to be optimum. A different proof, which has a weaker upper bound, appears in [17, Theorem 6], and related results are given in [6], [16], [23], and [23, Theorem 1].
5.2 FFT of Power Exponential Sum with Relatively Prime Index

### v14

# Simple Proof of the Primitive Root Conjecture

@@ Abstract
Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)\mid p-1$ be the multiplicative order of $u\operatorname{mod}p$. Define a prime counting function by $\pi(u,x)=\#\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}$. In 1967 Hooley proved a conditional asymptotic formula $\pi(u,x)=\delta(u)x(\log x)^{-1}+O(\log\log x(\log x)^{-2}$ for the primitive root conjecture. This note proves an unconditional asymptotic formula $\pi(u,x)=\delta(u)x(\log x)^{-1}+O(x(\log x)^{-2}$ of the same result, where $\delta(u)>0$ is the density constant.
@@ 1 Introduction
The symbols $\mathbb{N}=\{0,1,2,3,\ldots\}$, $\mathbb{Z}=\{0,\pm 1,\pm 2,\ldots\}$ and $\mathbb{P}=\{2,3,5,\ldots\}$ denote the set of natural numbers, the set integers and the set of prime numbers respectively. Let $x\geq 1$ be a large real number and let $\pi(x)=\#\{\text{prime }p\leq x\}$ be the prime counting function. The multiplicative order of an element in a finite field $u\in\mathbb{F}_{p}$ is denoted by $\operatorname{ord}_{p}u$, see Definition 2.1. The constant $\delta(u)\geq 0$ is the density
$\delta(u)=\lim_{x\to\infty}\frac{\#\left\{p\in\mathbb{P}:\operatorname{ord}_{p}(u)=p-1\right\}}{\pi(x)}$ (1.1)
of the subset of primes $\mathscr{P}_{u}=\left\{p\in\mathbb{P}:\operatorname{ord}_{p}(u)=p-1\right\}$ with a fixed primitive root $u$ in the set of primes $\mathbb{P}$. The expected number of primes $p\leq x$ with a fixed primitive root $u\operatorname{mod}p$ has the asymptotic formula
@@
A conditional proof of this result was achieved in [21], and simplified sketches of the proof appear in [29, p. 8], and similar references. The determination of the constant $\delta(u)\geq 0$ for a fixed integer $u\in\mathbb{Z}$ is an interesting technical problem, [21, p. 218], [26], [25], et alii. An introduction to its historical development, and its calculations is covered in [29, pp. 3–10], and [41]. More recent results are given in [43], [37], et alia.
@@
$\frac{1}{x}\sum_{u\leq x}\pi_{u}(x)=a_{0}\operatorname{li}(x)+O\left(\frac{x}{(\log x)^{-c}}\right).$ (1.3)
The Artin constant is defined by
$a_{0}=\sum_{n\geq 1}\frac{\mu(n}{[\mathcal{K}_{n}:\mathbb{Q}]}=\prod_{p\geq 2}\left(1-\frac{1}{p(p-1)}\right)$ (1.4)
where $\mathcal{K}_{n}:\mathbb{Q}]$ is a number field extension of $\mathbb{Q}$ and $c>1$ is an arbitrary number, was proved in [41]. These works had shown that almost all admissible integers $u$ are primitive roots for infinitely many primes. The number of exceptions is a subset of zero density in $\mathbb{Z}$. The individual quantity $\pi_{u}(x)$ in (1.2) can be slightly different from the average quantity in (1.4). The variations, discovered by the Lehmers using numerical experiments, depend on the primes decomposition of the fixed value $u$, see [21, p. 220], [29, p. 3], and similar references for the exact formula for the density $\delta(u)\geq 0$.
These results have been extended to quadratic numbers fields in [5], [NW2000], [38], and other number fields in [22], [1], [30], et cetera. Other related results are given in [31], [16], [36], [2], et alii. The Artin primitive root conjecture for functions fields was proved by Bilharz, see [39, Theorem 10.11], and the same conjecture for polynomials over finite fields was proved in [35]. The literature on this topic is vast and extensive.
This note introduces a simple unconditional proof of the primitive root conjecture over the integers $\mathbb{Z}$.
@@
where $\operatorname{li}(x)$ is the logarithm integral, and $\delta(u)>0$ is a constant, for all large numbers $x\geq 1$ unconditionally.
Section 2 to Section 6 introduce the notation and some standard results related to or applicable to the investigation of primitive roots in cyclic groups. Section 7 presents a proof of Theorem 1.1.
2 Characteristic Functions in Finite Fields
The characteristic function $\Psi:G\longrightarrow\{0,1\}$ of primitive elements is one of the standard analytic tools employed to investigate the various properties of primitive roots in cyclic groups $G$. Many equivalent representations of the characteristic function $\Psi$ of primitive elements are possible, a few are investigated here.
Definition 2.1.
The class function $\operatorname{ord}:\mathbb{F}_{p}^{\times}\longrightarrow\{d<p:d\mid p-1\}$ defined by $\operatorname{ord}_{p}u=\min\#\{n<p:u^{n}\equiv 1\bmod p\}$ measures the multiplicative order of an element $u\neq 0$ in the multiplicative group $\mathbb{F}_{p}^{\times}$. An element of maximal multiplicative order $\operatorname{ord}_{p}u=p-1$ is called a primitive element.
2.1 Divisor Dependent Characteristic Function
The divisor-dependent representation of the characteristic function detects the multiplicative order of an element by means of the divisors of the totient function $\varphi(p)=p-1$. The version for primitive elements in the prime finite fields $\mathbb{F}_{p}$ is described below.
@@ 2.1 Divisor Dependent Characteristic Function
Let $p\geq 2$ be an integer and let $\chi$ be a multiplicative character of order $\operatorname{ord}\chi=d$. If $u\in\mathbb{F}_{p}$ is a nonzero element, then
@@
where $\mu:\mathbb{N}\longrightarrow\{-1,0,1\}$ is the Mobius function.
There are a few other variant proofs of this result, these are widely available in the literature. Almost every result in the theory of primitive roots in finite fields is based on this characteristic function, but sometimes written in different forms. This technique was developed by Landau, Vinogradov, Davenport, Erdos, and other authors, see [24], [44], [6], [12], [47], et alii. An extension of this characteristic function to the finite ring $\mathbb{Z}/n\mathbb{Z}$ is presented in [28, Lemma 4]. The main obstacle in this technique is the primes decomposition of the totients p1p-1. Specifically, the finite sum $\sum_{d\mid p-1}|\mu(d)|\ll p^{\varepsilon}$ is an obstacle in many applications.
2.2 Divisorfree Characteristic Functions
Anew divisorfree representation of the characteristic function of primitive element is introduced in this section. It detects the order $\text{ord}_{p}(u)\geq 1$ of the element $u\in\mathbb{F}_{p}$ by means of the solutions of the equation $\tau^{s}-u=0$ in $\mathbb{F}_{p}$, where $u,\tau$ are constants, and $s\in\mathscr{R}=\{s<p:\gcd(s,p-1)=1\}$ is a variable.
Lemma 2.2.
Let $p\geq 2$ be a prime and let $\tau$ be a primitive root mod $p$ and let $\psi\neq 1$ be a nonprincipal additive character of order $\operatorname{ord}\psi=p$. If $u\in\mathbb{F}_{p}$ is a nonzero element, then
@@ 2.2 Divisorfree Characteristic Functions
Proof.
Set the additive character $\psi(s)=e^{i2\pi as/p}\in\mathbb{C}$. Now, as the index $s\in\mathscr{R}$ ranges over the integers relatively prime to $\varphi(p)=p-1$, the element $\tau^{n}\in\mathbb{F}_{p}^{\times}$ ranges over the primitive roots modulo $p$. Accordingly, the equation
$a=\tau^{n}-u=0$ (2.1)
has a unique solution $s\in\mathscr{R}$ if and only if the fixed element $u\in\mathbb{F}_{p}$ is a primitive root. This implies that the inner sum in
[…]
collapses to $\sum_{0\leq s<p}e^{i2\pi as/p}=\sum_{0\leq s<p}1=p$. Otherwise, if the element $u\in\mathbb{F}_{p}$ is not a primitive root, then the equation (2.1) has no solution $n\in\mathscr{R}$, and the inner sum in (2.2) collapses to $\sum_{0\leq s<p}e^{i2\pi as/p}=0$, this follows from the geometric series formula $\sum_{0\leq n\leq N-1}w^{n}=(w^{N}-1)/(w-1)$, where $w=e^{i2\pi a/(p-1)}\neq 1$ and $N=p$. This completes the verification. ∎
Another approach to the construction of divisorfree characteristic function is via the discrete logarithm $\log_{\tau}:\mathbb{F}_{p}^{\times}\longrightarrow\mathbb{F}_{p}$ in finite field $\mathbb{F}_{p}$.
@@
Let $p\geq 2$ be a prime and let $\tau$ be a primitive root mod $p$ and let $\psi\neq 1$ be a nonprincipal additive character of order $\operatorname{ord}\psi=p$. If $u\in\mathbb{F}_{p}$ is a nonzero element, then
@@
Proof.
Let $\mathscr{R}=\{s<p:\gcd(s,p-1)=1\}$ and set the additive character $\psi(s)=e^{i2\pi as/p}\in\mathbb{C}$. Consider a fixed element $u=\tau^{s}\in\mathbb{F}_{p}$ and its discrete logarithm $\log_{\tau}u=s$. As the index varies over the set $s\in\mathscr{R}$ ranges over the integers relatively prime to $\varphi(p)=p-1$, the equation
$a=s-\log_{\tau}u=0$ (2.3)
has a unique solution $s=\log_{\tau}u$ if and only if the fixed element $u\in\mathbb{F}_{p}$ is a primitive root. This implies that the inner sum in
@@
collapses to $\sum_{0\leq s<p}e^{i2\pi as/p}=\sum_{0\leq s<p-1}1=p$. Otherwise, if the element $u-1\in\mathbb{F}_{p}$ is not a primitive root, then the equation (2.3) has no solution $s\in\mathscr{R}$ , and the inner sum in (2.4) collapses to $\sum_{0\leq s<p-1}e^{i2\pi as/p}=0$, this follows from the geometric series formula $\sum_{0\leq n\leq N-1}w^{n}=(w^{N}-1)/(w-1)$, where $w=e^{i2\pi a/p}\neq 1$ and $N=p$. This completes the verification. ∎
@@ 4 Estimates of Power Exponential Sums
where $\omega=e^{i2\pi/p}$ and $x\leq p-1$. The case of interest here is the exponential sum
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}.$ (4.2)
Substituting $t=0$ in (4.1) contributes $\varphi(p-1)/p$, and rearranging it yield
$=$ $\sum_{1\leq n\leq p-1}\frac{1}{p}\sum_{1\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d)-\frac{\varphi(p-1)}{p}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}1\leq n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p-1)}{p}.$
@@
There are many proofs of this result, consult [32], [15, Theorem 6], and related results are given in [14], [19], [4], and [20, Theorem 1]. The upper bound given in Theorem 4.1 seems to be optimum. A different proof, which has a weaker upper bound, appears in [15, Theorem 6], and related results are given in [4], [13], [20], and [20, Theorem 1].∎
4.1 Finite Fourier Transform of the Power Exponential Sum